# Congruences between Klingen-Eisenstein series and cusp forms on $\mathrm{U}_{n,n}$

### 09 Mar 2026 v1

@@ Abstract.
Hermitian modular forms, differential operators, pullback formula
@@ 1. Introduction
A foundational instance of this principle is the work of Skinner and Urban [32], who established the Main Conjecture for $\mathrm{GL}_{2}$. Their strategy involved constructing congruences between Eisenstein series and cusp forms on the unitary group $\mathrm{GU}(2,2)$.
@@
Specifically, let $r<n$ be a positive integer. We consider a Hecke cusp form $f$ on the smaller unitary group $\mathrm{U}_{r,r}$ and its associated Klingen-Eisenstein series $f]_{r}^{n}$ on $\mathrm{U}_{n,n}$. The main result of this paper (Theorem 6.2) asserts that if the algebraic part of the special value of the standard $L$-function associated with $f$, denoted by $L(s,f,\mathrm{St})$, is divisible by a prime ideal $\mathfrak{p}$ (under appropriate normalization and conditions on the prime $p$), then there exists a Hecke cusp form $F$ on $\mathrm{U}_{n,n}$ such that
@@
where $\mathfrak{P}$ is a prime ideal of the Hecke field lying above $\mathfrak{p}$, and $\equiv_{ev}$ denotes the congruence of eigenvalues for all Hecke operators in the spherical Hecke algebra.
To prove this theorem, we first utilize the differential operators $\mathbb{D}_{\bm{k},\bm{l}}$ constructed in our previous work [34], which are designed to ensure that the pullback of an automorphic form remains automorphic on the subgroup. By applying these operators to the Eisenstein series on $\mathrm{U}_{n,n}$ and restricting it to $\mathrm{U}_{n_{1},n_{1}}\times\mathrm{U}_{n_{2},n_{2}}$, we derive a pullback formula. At the same time, we establish the necessary arithmetic framework by defining an integral structure on the space of automorphic forms and proving the rationality of the space of Hermitian automorphic forms.
@@
(See §5 for the precise definitions and normalization.)
@@
where $v_{\mathfrak{p}}$ denotes the $\mathfrak{p}$-adic valuation, and $A_{[f]_{\mu}^{n}}(\ast,S_{0})$ is an $S_{0}$-th Fourier coefficient
of $f]_{\mu}^{n}$. We further impose several natural technical conditions on $\mathfrak{p}$ ensuring integrality (see Theorem 6.2).
@@
This paper is organized as follows:
Section 2 provides the preliminary definitions of Hermitian automorphic forms on $\mathrm{U}_{n,n}$ and their Fourier expansions.
Section 3 is devoted to the local theory of Hecke algebras. We describe the structure of the spherical Hecke algebra at inert, ramified, and split places, and explicitly compute the action of Hecke operators on the Fourier coefficients. This explicit description is crucial for establishing the integrality of the Hecke action.
Section 4 introduces the differential operators and the Hermitian Eisenstein series. We then state the pullback formula (Theorem 4.8).
Section 5 discusses the arithmetic properties of Hermitian automorphic forms. We prove the rationality of the space of automorphic forms (Theorem 5.12) and establish the integrality properties of Klingen-Eisenstein series.
Section 6 contains the main result. We combine the results of the previous sections to prove the existence of congruences between Klingen-Eisenstein series and cusp forms (Theorem 6.2), provided the algebraic part of the standard $L$-value satisfies the divisibility condition.
Section 7 provides explicit examples of congruences between Klingen-Eisenstein series and cusp forms.
Acknowledgment
The author would like to thank T. Ikeda for his great guidance and support as the author’s doctoral advisor, and H. Katsurada for his many suggestions and comments.
This work was supported by JST SPRING, Grant Number JPMJSP2110.
@@ 2. Hermitian automorphic forms
We put $m=m_{F}:=\#\mathbf{a}=[F:\mathbb{Q}]$. We put $K_{v}=\prod_{w|v}K_{w}$ and $\mathcal{O}_{K_{v}}=\prod_{w|v}\mathcal{O}_{K_{w}}$ for a place $v$ of $F$. Let $\mathbb{A}=\mathbb{A}_{F}$ be the adele ring of $F$, and $\mathbb{A}_{0}$, $\mathbb{A}_{\infty}$ the finite and infinite parts of $\mathbb{A}$, respectively.
@@ 4.3. Pullback formula
Let $n_{1}$ and $n_{2}$ be positive integers with $n_{1}\geq n_{2}$, and put $n=n_{1}+n_{2}$. Let $m=m_{F}$ be the number of infinite places of $F$. Let $\kappa=(\kappa_{v})_{v},\nu=(\nu_{v})_{v}$ be families of positive integers, such that $\kappa_{v}+\nu_{v}=\mu$ for each infinite place $v$ of $F$, where $\mu$ is the fixed positive integer as above. and let $\bm{k}=(\bm{k}_{v})_{v}$ and $\bm{l}=(\bm{l}_{v})_{v}$ be families of dominant integral weights such that $\ell(\bm{k}_{v})<n_{2}$, $\ell(\bm{l}_{v})\leq n_{2}$, and $\ell(\bm{k}_{v})+\ell(\bm{l}_{v})\leq\kappa_{v}+\nu_{v}$ for each infinite place $v$ of $F$. We put $\rho^{\prime}_{n_{1}}={\det}^{\kappa}\rho_{n_{1},\bm{l}}\boxtimes{\det}^{\nu}\rho_{n_{1},\bm{k}}$ and $\rho_{n_{2}}={\det}^{\kappa}\rho_{n_{2},\bm{k}}\boxtimes{\det}^{\nu}\rho_{n_{2},\bm{l}}$, and fix the differential operator $\mathbb{D}_{\bm{k},\bm{l}}=\mathbb{D}_{\bm{k},\bm{l}}^{(n_{1},n_{2})}=P_{\bm{k},\bm{l}}(\pi^{+})$ satisfying condition (A) for ${\det}^{\kappa}\boxtimes{\det}^{\nu}$ and $\rho^{\prime}_{n_{1}}\boxtimes\rho_{n_{2}}$, as defined in [33, Theorem 5.13]. Hereafter, $\mathbb{D}_{\bm{k},\bm{l}}$ denotes this uniquely determined differential operator.
@@
Assume that $\mathrm{Re}(s)>n$. Let the above notation and assumptions stand. Then, for any Hecke cusp form $f\in\mathcal{S}^{\sharp}_{n_{2}}(\rho_{n_{2}})$, we have
$\left(f,\,(\mathbb{D}_{\bm{k},\bm{l}}E_{n,\kappa})(\iota_{n_{1},n_{2}}(g_{1},*),\overline{s};\chi)\right)=c(s,\rho_{n_{2}})\,D(s,f;\overline{\chi})\,\overline{[f^{\dagger}_{\chi}]_{\kappa}^{n_{1}}(g_{1},s;\chi)},$
where $c(s,\rho_{n_{2}})$ is a function depending only on $\rho_{n_{2}}$, which is defined in [34].
@@
The constant $c(s,\rho_{n_{2}})$ in the theorem is generally not explicitly stated, but in certain cases its value can be computed. We set $|\bm{k}|=\sum_{v,i}k_{v,i}$ and $|\bm{l}|=\sum_{v,i}l_{v,i}$ for the tuple of dominant integral weights $(\bm{k},\bm{l})$.
@@
Let the above notions and assumptions stand. Assume that $\kappa_{v}+\nu_{v}=\mu$ for all $v\in\mathbf{a}$. Then the constant $c(\mu/2,\rho_{n_{2}})$ appearing in Theorem 4.8 is given by
$c(\mu/2,\rho_{n_{2}})={}$ $2^{-2mn_{2}^{2}-m(n_{2}-2)\mu+|\bm{k}|+|\bm{l}|}\,\pi^{mn_{2}^{2}}$
@@ 5. Integrality of Hermitian automorphic forms
$c_{r}=(2\pi\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}\,c(\mu/2,\rho_{r})\cdot\frac{D_{K}^{-r/2}\prod_{i=0}^{2r-1}L_{F}(1-\mu+i,\chi_{-D_{K}}^{i})\,L_{F}(\mu-i,\chi_{-D_{K}}^{i})^{-1}}{\prod_{j=1}^{r}\Gamma_{\mathbb{C}}(\mu+k_{j}-j-r+1)\,\Gamma_{\mathbb{C}}(\mu+l_{j}-j)}.$
@@
$(\sqrt{-1})^{|\rho_{r}|}\,\epsilon_{\mu}(\gamma_{(n_{1})},\gamma_{(n_{2})};S_{1},S_{2})\in(V_{\rho_{n_{1}}}\otimes V_{\rho^{\prime}_{n_{2}}})(\mathbb{Q})$
@@
$(\sqrt{-1})^{|\rho_{r}|}\,\epsilon_{\mu}(\gamma_{(n_{1})},\gamma_{(n_{2})};S_{1},S_{2})\in(V_{\rho_{n_{1}}}\otimes V_{\rho^{\prime}_{n_{2}}})(\mathbb{Z}_{(p)}),$
@@
$2^{-(|\bm{k}|+|\bm{l}|)}\,\mathbb{D}_{\bm{k},\bm{l}}f(\iota)=(P_{\bm{k},\bm{l}}(\partial_{Z})f_{1},\ldots,P_{\bm{k},\bm{l}}(\partial_{Z})f_{h})^{\sharp},$
@@
Let $\mu$ be a positive even integer with $\mu\geq 2n$. If $F$ is a Hecke eigenform in $\mathcal{A}^{\sharp}_{n}(\rho_{n})$, then there exists a totally real algebraic field $L$ such that
@@
When $n=1$, we have
$L\bigl(s_{0}+\tfrac{1}{2},\,f,\,\mathrm{St}\bigr)=L\bigl(s+\tfrac{k}{2},f\bigr)\,L\bigl(s+\tfrac{k}{2},f\otimes\chi_{-D_{K}}\bigr),$
where $f$ is a Hecke cusp form in $\mathcal{S}^{\sharp}_{1}({\det}^{k/2}\boxtimes{\det}^{k/2})=S_{k}(\mathrm{SL}_{2}(\mathbb{Z}))$ and $L(s,f)$ denotes the classical $L$-function of $f$, that is,
$L(s,f)=\sum_{n=1}^{\infty}a(f,n)n^{-s}=\prod_{p}(1-a(f,p)p^{-s}+p^{k-1-2s})^{-1}.$
Moreover, the Petersson inner product $(f,f)$ defined on the Hermitian upper half-plane differs from the classical Petersson inner product on the Siegel upper half-plane by a rational factor (see, for example, the appendix to Section 10 of Ichino–Ikeda [17]). Therefore, the above identity is compatible with Shimura [28].
@@ 7. Example
We identify the representation space of $\det^{4}\mathrm{Sym}^{7}\boxtimes\det^{4}\mathrm{Sym}^{7}$ with the space of homogeneous polynomials of degree $7$ in the variables $u_{1},u_{2}$ and of degree $7$ in the variables $v_{1},v_{2}$. Since $\dim S_{22}(\Gamma_{K}^{(1)})=1$, we may normalize the factor so that $\mathfrak{A}(\Delta_{22})=1$. By direct computation, we have
@@
$=$ $\frac{1}{D}\bigg[-3\cdot 107\cdot 7333\cdot 9623\left(u_{1}^{7}v_{1}^{7}+u_{2}^{7}v_{2}^{7}\right)$
@@
where $D=2^{9}\cdot 3^{7}\cdot 5^{3}\cdot 7^{3}\cdot 11\cdot 13\cdot 52140059$.
@@
Since these primes satisfy the hypotheses of Theorem 6.2, we conclude that for each such $p$, there exists a Hecke cusp form $F\in\mathcal{S}^{\sharp}_{2}\bigl((\det^{14}\mathrm{Sym}^{8})\boxtimes 1\bigr)$ such that
@@
From the above computations together with some additional straightforward verifications, we conclude that $809$ satisfies the assumptions of Corollary 6.5. Hence there exists a Hecke eigenform $F\in\mathcal{A}^{\sharp}_{3}(\rho_{3})$ such that

### 20 Mar 2026 v2

@@ Abstract.
Hermitian modular form, Unitary groups, Congruences
@@ 1. Introduction
A foundational instance of this principle is the work of Skinner and Urban [32], who established the Main Conjecture for $\mathrm{GL}_{2}$. Their strategy involved constructing congruences between Eisenstein series and cusp forms on the unitary similitude group $\mathrm{GU}(2,2)$.
@@
Specifically, let $r<n$ be a positive integer. We consider a Hecke cusp form $f$ on the smaller unitary group $\mathrm{U}_{r,r}$ and its associated Klingen-Eisenstein series $f]_{r}^{n}$ (defined in Section 4.2) on $\mathrm{U}_{n,n}$. The main result of this paper (Theorem 6.2) asserts that if the algebraic part of the special value of the standard $L$-function associated with $f$, denoted by $L(s,f,\mathrm{St})$, is divisible by a prime ideal $\mathfrak{p}$ (under appropriate normalization and conditions on the prime $p$), then there exists a Hecke cusp form $F$ on $\mathrm{U}_{n,n}$ such that
@@
where $\mathfrak{P}$ is a prime ideal of the Hecke field lying above $\mathfrak{p}$, and $\equiv_{ev}$ denotes the congruence of eigenvalues for all Hecke operators in the spherical Hecke algebra discussed in Section 3.
To prove this theorem, we first utilize the vector-valued differential operators $\mathbb{D}_{\bm{k},\bm{l}}$ constructed in our previous work [34], which are designed to ensure that the pullback of an automorphic form remains automorphic on the subgroup. By applying these operators to the Eisenstein series on $\mathrm{U}_{n,n}$ and restricting it to $\mathrm{U}_{n_{1},n_{1}}\times\mathrm{U}_{n_{2},n_{2}}$, we derive a pullback formula (Theorem 4.8). At the same time, we establish the necessary arithmetic framework by defining an integral structure on the space of automorphic forms and proving the rationality of the space of Hermitian automorphic forms (Theorem 5.12).
@@
(See Section 5 for the precise definitions and normalization.)
@@
where $v_{\mathfrak{p}}$ denotes the $\mathfrak{p}$-adic valuation, and $A_{[f]_{\mu}^{n}}(\ast,S_{0})$ is an $S_{0}$-th Fourier coefficient of $f]_{\mu}^{n}$ defined as (2.6). We further impose several natural technical conditions on $\mathfrak{p}$ ensuring integrality.
@@
This paper is organized as follows: Section 2 provides the preliminary definitions of Hermitian automorphic forms on $\mathrm{U}_{n,n}$ and their Fourier expansions. Section 3 is devoted to the local theory of Hecke algebras, describing the structure at inert, ramified, and split places. Section 4 introduces the differential operators and states the pullback formula (Theorem 4.8). Section 5 discusses the arithmetic properties and rationality of the space of automorphic forms (Theorem 5.12). Section 6 contains the proof of the main result. Section 7 provides explicit numerical examples for $n=2$ and $n=3$.
Acknowledgements
The author is grateful to T. Ikeda for his guidance and support as the author’s doctoral advisor, and to H. Katsurada for valuable comments. This work was supported by the Japan Science and Technology Agency (JST) SPRING Program, Grant Number JPMJSP2110.
@@ 2. Hermitian automorphic forms
We put $m=m_{F}:=\#\mathbf{a}=[F:\mathbb{Q}]$. We put $K_{v}=\prod_{w|v}K_{w}$ and $\mathcal{O}_{K_{v}}=\prod_{w|v}\mathcal{O}_{K_{w}}$ for a place $v$ of $F$. Let $\mathbb{A}=\mathbb{A}{F}$ be the adele ring of $F$, and $\mathbb{A}_{0}$, $\mathbb{A}_{\infty}$ the finite and infinite parts of $\mathbb{A}$, respectively.
@@ 4.3. Pullback formula
Let $n_{1}$ and $n_{2}$ be positive integers with $n_{1}\geq n_{2}$, and put $n=n_{1}+n_{2}$. Let $m=m_{F}$ be the number of infinite places of $F$. Let $\kappa=(\kappa_{v})_{v},\nu=(\nu_{v})_{v}$ be families of positive integers, such that $\kappa_{v}+\nu_{v}=\mu$ for each infinite place $v$ of $F$, where $\mu$ is the fixed positive integer as above. and let $\bm{k}=(\bm{k}_{v})_{v}$ and $\bm{l}=(\bm{l}_{v})_{v}$ be families of dominant integral weights such that $\ell(\bm{k}_{v})<n_{2}$, $\ell(\bm{l}_{v})\leq n_{2}$, and $\ell(\bm{k}_{v})+\ell(\bm{l}_{v})\leq\kappa_{v}+\nu_{v}$ for each infinite place $v$ of $F$. We put $\rho_{n_{1}}={\det}^{\kappa}\rho_{n_{1},\bm{k}}\boxtimes{\det}^{\nu}\rho_{n_{1},\bm{l}}$ and $\rho^{\prime}_{n_{2}}={\det}^{\kappa}\rho_{n_{2},\bm{l}}\boxtimes{\det}^{\nu}\rho_{n_{2},\bm{k}}$, and fix the differential operator $\mathbb{D}_{\bm{k},\bm{l}}=\mathbb{D}_{\bm{k},\bm{l}}^{(n_{1},n_{2})}=P_{\bm{k},\bm{l}}(\pi^{+})$ satisfying condition (A) for ${\det}^{\kappa}\boxtimes{\det}^{\nu}$ and $\rho_{n_{1}}\boxtimes\rho^{\prime}_{n_{2}}$, as defined in [33, Theorem 5.13]. Hereafter, $\mathbb{D}_{\bm{k},\bm{l}}$ denotes this uniquely determined differential operator.
@@
Assume that $\mathrm{Re}(s)>n$. Let the above notation and assumptions stand. Then, for any Hecke cusp form $f\in\mathcal{S}^{\sharp}_{n_{2}}(\rho^{\prime}_{n_{2}})$, we have
$\left(f,\,(\mathbb{D}_{\bm{k},\bm{l}}E_{n,\kappa})(\iota_{n_{1},n_{2}}(g_{1},*),\overline{s};\chi)\right)=c(s,\rho^{\prime}_{n_{2}})\,D(s,f;\overline{\chi})\,\overline{[f^{\dagger}_{\chi}]_{\kappa}^{n_{1}}(g_{1},s;\chi)},$
where $c(s,\rho^{\prime}_{n_{2}})$ is a function depending only on $\rho^{\prime}_{n_{2}}$, which is defined in [34].
@@
The constant $c(s,\rho^{\prime}_{n_{2}})$ in the theorem is generally not explicitly stated, but in certain cases its value can be computed. We set $|\bm{k}|=\sum_{v,i}k_{v,i}$ and $|\bm{l}|=\sum_{v,i}l_{v,i}$ for the tuple of dominant integral weights $(\bm{k},\bm{l})$.
@@
Let the above notions and assumptions stand. Assume that $\kappa_{v}+\nu_{v}=\mu$ for all $v\in\mathbf{a}$. Then the constant $c(\mu/2,\rho^{\prime}_{n_{2}})$ appearing in Theorem 4.8 is given by
$c(\mu/2,\rho^{\prime}_{n_{2}})={}$ $2^{-2mn_{2}^{2}-m(n_{2}-2)\mu+|\bm{k}|+|\bm{l}|}\,\pi^{mn_{2}^{2}}$
@@ 5. Integrality of Hermitian automorphic forms
$c_{r}=(2\pi\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}\,c(\mu/2,\rho^{\prime}_{r})\cdot\frac{D_{K}^{-r/2}\prod_{i=0}^{2r-1}L_{F}(1-\mu+i,\chi_{-D_{K}}^{i})\,L_{F}(\mu-i,\chi_{-D_{K}}^{i})^{-1}}{\prod_{j=1}^{r}\Gamma_{\mathbb{C}}(\mu+k_{j}-j-r+1)\,\Gamma_{\mathbb{C}}(\mu+l_{j}-j)}.$
@@
$(\sqrt{-1})^{|\rho_{n_{2}}|}\,\epsilon_{\mu}(\gamma_{(n_{1})},\gamma_{(n_{2})};S_{1},S_{2})\in(V_{\rho_{n_{1}}}\otimes V_{\rho^{\prime}_{n_{2}}})(\mathbb{Q})$
@@
$(\sqrt{-1})^{|\rho_{n_{2}}|}\,\epsilon_{\mu}(\gamma_{(n_{1})},\gamma_{(n_{2})};S_{1},S_{2})\in(V_{\rho_{n_{1}}}\otimes V_{\rho^{\prime}_{n_{2}}})(\mathbb{Z}_{(p)}),$
@@
$2^{-(|\bm{k}|+|\bm{l}|)}\,\mathbb{D}_{\bm{k},\bm{l}}f=(P_{\bm{k},\bm{l}}(\partial_{Z})f_{1},\ldots,P_{\bm{k},\bm{l}}(\partial_{Z})f_{h})^{\sharp},$
@@
Let $\mu$ be a positive even integer with $\mu\geq 2n$. If $F$ is a Hecke eigenform in $\mathcal{A}^{\sharp}_{n}(\rho_{n})$, then
@@
When $n=1$ and the class number of $K$ is $1$, we have
$L\bigl(s_{0}+\tfrac{1}{2},\,f^{\sharp},\,\mathrm{St}\bigr)=L\bigl(s_{0}+\tfrac{k}{2},\,f\bigr)L\bigl(s_{0}+\tfrac{k}{2},\,f\otimes\chi_{K}\bigr),$
where $f\in S_{k}(\mathrm{SL}_{2}(\mathbb{Z}))$ is a classical Hecke cusp form and $L(s,f)$ denotes its Dirichlet $L$-series:
$L(s,f)=\sum_{n=1}^{\infty}a(n,f)n^{-s}=\prod_{p}(1-a(p,f)p^{-s}+p^{k-1-2s})^{-1}.$
Moreover, the Petersson norm $(f^{\sharp},f^{\sharp})$ on the Hermitian upper half-plane coincides with the classical one (cf. [17, Appendix to §10]). Thus, our result is consistent with the period relations shown by Shimura [28].
@@ 7. Example
We identify the representation space of $\det^{4}\mathrm{Sym}^{7}\boxtimes\det^{4}\mathrm{Sym}^{7}$ with the space of homogeneous polynomials of degree $7$ in the variables $u_{1},u_{2}$ and of degree $7$ in the variables $v_{1},v_{2}$. Since $\dim\mathcal{S}_{1}(\det^{11}\boxtimes\det^{11})=1$, we have $\mathfrak{A}(\Delta_{22})=1$. By direct computation, we have
@@
$=$ $\frac{1}{N}\bigg[-3\cdot 107\cdot 7333\cdot 9623\left(u_{1}^{7}v_{1}^{7}+u_{2}^{7}v_{2}^{7}\right)$
@@
where $N=2^{9}\cdot 3^{7}\cdot 5^{3}\cdot 7^{3}\cdot 11\cdot 13\cdot 52140059$.
@@
Since these primes satisfy the hypotheses of Theorem 6.2, we conclude that for each such $p$, there exists a Hecke cusp form $F\in\mathcal{S}^{\sharp}_{2}\bigl(\det^{4}\mathrm{Sym}^{7}\boxtimes\det^{4}\mathrm{Sym}^{7}\bigr)$ such that
@@
From the above computations together with some additional straightforward verifications, we conclude that $809$ satisfies the assumptions of Corollary 6.5. Hence there exists a Hecke eigenform $F\in\mathcal{A}^{\sharp}_{3}(\det^{6}\boxtimes\det^{6})$ such that